OpenStudy (anonymous):

whats the answer to 12 over one when simplify?

myininaya (myininaya):

a/1=a

myininaya (myininaya):

\[\frac{ :)}{1}=:)\]

jimthompson5910 (jim_thompson5910):

Ie any number divided by 1 is that same number

jimthompson5910 (jim_thompson5910):

so 12/1 = 12

myininaya (myininaya):

\[\frac{can of tuna fish}{1}=can of tuna fish\]